Determine the Volume of a Solid by Integrating a Cross-Section With a Circle or Semicircle Question The base formed by slicing through the center of a solid S is the ellipse + y,-1. The cross sections pe the base and the x-axis are circles. Find the volume of S. Enter your answer in terms of r. 64 9 Provide your answer below: MODE INSTRUCTION MI
Answer #1
  1. x^2/64+y^2/9=1y 2 + il Put y= 0 25 +0=1 1 2=64 x = 8 + V -8 <x < 8 and and hey are the one y2 = 9 11-es) - 0 lw=alradius) Area of a circle